# language=R,texcl=true,style=vs

### v1

@@ 2.1 Energy distance
2.2 Testing for equality of two univariate distributions
To test for the equality of two univariate and multivariate distributions, the test statistic is the energy distance (1). Since we focus on the univariate case and assume that we have two vectors of samples ${\bf x}=(x_{1},\ldots,x_{n})$ and ${\bf y}=(y_{1},\ldots,y_{m})$, Eq. (1) is written in its univariate form as
$\mathcal{\varepsilon}\left({\bf x},{\bf y}\right)=\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}|x_{i}-y_{j}|-\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}|x_{i}-x_{j}|-\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}|y_{i}-y_{j}|,$ (2)
and the test statistic now becomes $T=\frac{nm}{n+m}\mathcal{\varepsilon}\left({\bf x},{\bf y}\right)$.
Permutations are used, where shuffling takes place between the two vectors and Eq. (2) is computed. This process is repeated $R$ times, and the proportion of times the energy distance test statistics of the permuted data exceeds the energy distance test statistic computed at the observed data serves as the approximate p-value.
@@ 2.4.1 Computation of distance variance, covariance, correlation and partial correlation
where $x_{(i)}$ denote the ordered values, in ascending order. Regarding the denominator of $\gamma_{1}$, we can dot the following
$\sum_{i\neq j}^{n}a_{ij}^{2}=\sum_{i\neq j}^{n}|x_{i}-x_{j}|=n\sum_{i=1}^{n}x^{2}-\left(\sum_{i=1}^{n}x\right)^{2}.$
@@ 2.4.2 Computation of energy distance
2.4.3 Computations for testing the equality of two univariate distributions
It was mentioned earlier that random permutations occur repeatedly. This means that the constant factor $\frac{nm}{n+m}$ that multiplies the energy distance $(\varepsilon({\bf x},{\bf y}))$ does not influence the p-value. Let us now break Eq. (2) into its three components
@@ 2.4.3 Computations for testing the equality of two univariate distributions
The first component ($\gamma_{1}$) never changes during random permutations, therefore it does not need to be computed. However, the values of $\gamma_{2}$ and $\gamma_{3}$ change. Fortunately, in the univariate case, we can again use the identity used in the Gini coefficient to compute the sum of all pairwise distances, without actually computing the distances
@@
The same formula applies to the $\gamma_{3}$ quantity that refers to the second vector. Some extra tricks take place, for instance minimization of the repeated computations such as sums and the creation of the necessary sequences, and use of fast functions to sample the permutations and sort the vectors [^3] An R implementation of this test is given in the Appendix..
@@ 5 Conclusions
The R version of the command eqdist.etest() to test the equality of two univariate distributions
eqdist.etest <- function(y, x, R = 999) {
  #dii is the sum of all pairwise distances of x
  ni <- length(x)
  i <- 1:ni
  x <- Rfast::Sort(x)
  sx <- sum(x)
  dii <- 2 * sum(i * x ) - (ni + 1) * sx
  #djj is the sum of all pairwise distances of y
  nj <- length(y)
  j <- 1:nj
  y <- Rfast::Sort(y)
  sy <- sum(y)
  djj <- 2 * sum(j * y ) - (nj + 1) * sy
  s <- sx + sy  ## total sum
  n <- ni + nj  ## total sample size
  dis <- ( nj * dii / ni + ni * djj / nj ) / n
  ## Below are the permutations
  z <- c(x, y)
  pdis <- numeric(R)
  for ( k in 1:R ) {
    id <- Rfast2::Sample.int(n, n)
    z <- z[id]
    xp <- Rfast::Sort(z[i])
    sxp <- sum(xp)
    pdii <- 2 * sum(i * xp ) - (ni + 1) * sxp
    yp <- Rfast::Sort(z[-i])
    syp <- s - sxp
    pdjj <- 2 * sum(j * yp ) - (nj + 1) * syp
    pdis[k] <- ( nj * pdii / ni + ni * pdjj / nj ) / n
  ( sum(pdis >= dis) + 1 ) / (R + 1)
pdcor <- function(x, y, z) {
  a1 <- Rfast::dcor(x, y, bc = TRUE)\$dcor
  a2 <- Rfast::dcor(x, z, bc = TRUE)\$dcor
  a3 <- Rfast::dcor(y, z, bc = TRUE)\$dcor
  up <- a1 - a2 * a3
  down <- sqrt( 1 - a2^2 ) * sqrt( 1 - a3^2 )
  up / down

### v2

@@ 2.1 Energy distance
2.2 Testing for equality of two distributions
To test for the equality of two univariate and multivariate distributions, the test statistic is the energy distance (1) and the test statistic becomes $T=\frac{nm}{n+m}\mathcal{\varepsilon}\left({\bf x},{\bf y}\right)$. Permutations are used, where shuffling takes place between the two vectors and Eq. (LABEL:uenergy) is computed. This process is repeated $R$ times, and the proportion of times the energy distance test statistics of the permuted data exceeds the energy distance test statistic computed at the observed data serves as the approximate p-value.
@@ 2.4.1 Computation of distance variance, covariance, correlation and partial correlation
where $x_{(i)}$ denote the ordered values, in ascending order. Regarding the denominator of $\gamma_{1}$, we can do the following
$\sum_{i\neq j}^{n}a_{ij}^{2}=\sum_{i\neq j}^{n}|x_{i}-x_{j}|^{2}=n\sum_{i=1}^{n}x^{2}-\left(\sum_{i=1}^{n}x\right)^{2}.$
@@ 2.4.2 Computation of energy distance
2.4.3 Computations for testing the equality of two distributions
It was mentioned earlier that random permutations occur repeatedly. This means that the constant factor $\frac{nm}{n+m}$ that multiplies the energy distance $(\varepsilon({\bf x},{\bf y}))$ does not influence the p-value. In the case of univariate distributions we can break Eq. (1) into its three terms
@@ 2.4.3 Computations for testing the equality of two distributions
Fortunately, in the univariate case, we can again use the identity used in the Gini coefficient to compute the sum of all pairwise distances, without actually computing the distances
@@
The same formula applies to the $\gamma_{3}$ quantity that refers to the second vector. Regarding $\gamma_{1}$ this can be computed via the following identity
$\gamma_{1}=2\sum_{i=1}^{n+m}(iz_{(i)})-(n+m+1)\sum_{i=1}^{n+m}z_{i}-\gamma_{2}-\gamma_{3},$
where $z_{(1)},\ldots,z_{(n+m)}$ denotes the ordered observations of the combined samples, ${\bf z}=\left(x_{1},\ldots,x_{n},y_{1},\ldots,y_{m}\right)$. Some extra tricks take place, for instance, minimization of the repeated computations such as sums and the creation of the necessary sequences, and the use of fast functions to sample the permutations and sort the vectors.
For the case of multivariate distributions we can compute the first term ($\gamma_{1}$) and the other two ($\gamma_{2}$ and $\gamma_{3}$) using the commands dista(x, result = "sum") and Dist(x, result = "sum"), respectively, both available in the Rfast library. The resulting function [^3] See Appendix for an R implementation. is not faster than the available implementation in the energy library, but it is memory-saving. Note that the command Dist(x, result = "sum") can also facilitate the memory-efficient computation of the multivariate normality test (Székely and Rizzo,, 2023) as well.
@@ 5 Conclusions
pdcor <- function(x, y, z) {
  a1 <- Rfast::dcor(x, y, bc = TRUE)\$dcor
  a2 <- Rfast::dcor(x, z, bc = TRUE)\$dcor
  a3 <- Rfast::dcor(y, z, bc = TRUE)\$dcor
  up <- a1 - a2 * a3
  down <- sqrt( 1 - a2^2 ) * sqrt( 1 - a3^2 )
  up / down
The R version of the command eqdist.etest() to test the equality of two univariate or multivariate distributions
eqdist.etest <- function(y, x, R = 999) {
  if ( !is.matrix(x) ) {
    ## univariate
    #dii is the sum of all pairwise distances of x
    ni <- length(x)
    i <- 1:ni
    x <- Rfast::Sort(x)
    sx <- sum(x)
    dii <- 2 * sum(i * x) - (ni + 1) * sx
    #djj is the sum of all pairwise distances of y
    nj <- length(y)
    j <- 1:nj
    y <- Rfast::Sort(y)
    sy <- sum(y)
    s <- sx + sy
    djj <- 2 * sum(j * y) - (nj + 1) * sy
    n <- ni + nj  ## total sample size
    z <- Rfast::Sort( c(x, y) )
    ##dij <- Rfast::dista(x, y, result = "sum")
    dtot <- 2 * sum( ( 1:n ) * z ) - (n + 1) * s
    dij <- dtot - dii - djj
    stat <- dij - nj * dii / ni - ni * djj / nj
    pstat <- numeric(R)
    for ( k in 1:R ) {
      id <- Rfast2::Sample.int(n, ni)
      xp <- Rfast::Sort(z[id])
      sxp <- sum(xp)
      pdii <- 2 * sum(i * xp) - (ni + 1) * sxp
      yp <- Rfast::Sort(z[-id])
      syp <- s - sxp
      pdjj <- 2 * sum(j * yp) - (nj + 1) * syp
      pdij <- dtot - pdii - pdjj
      pstat[k] <- pdij - nj * pdii / ni - ni * pdjj / nj
    ## multivariate
  } else {
    nx <- dim(x)[1]  ;  ny <- dim(y)[1]
    n <- nx + ny
    stat <- Rfast::edist(x, y)
    z <- rbind(x, y)
    for ( k in 1:R ) {
      id <- Rfast2::Sample.int(n, nx)
      pstat[k] <- Rfast::edist(z[id, ], z[-id, ])
  ( sum( pstat >= stat ) + 1 ) / (R + 1)